The combined ages of Allison and Austin is 34. If Austin is 6 years older than Allison, how old is Austin?
Answer by ikleyn(52787)

.

Let x be the Austin's age.  Then Allison's age is (x-6) years.


Allison + Austin = 34


(x-6)   + x      = 34


2x = 34 + 6 = 40.


x = 20.


Answer.  Austin is 20 years old.

Solved.